OpenStudy (anonymous):

what is 2 times the square root of x,plus 2 times the square root of y ?

OpenStudy (anonymous):

\[2\sqrt{x}+2\sqrt{y}\]

OpenStudy (anonymous):

\[2(\sqrt{x}+\sqrt{y})\]